Steam, at 200° C is passed into a container of negligible heat capacity containing 40 g of ice and 200 g of water at 0° C until the ice is completely melted. Determine the total mass of water in the container
1
Expert's answer
2019-04-01T10:30:26-0400

Heat required by ice and water to go up to 100°C = m1L1 + m1c1 Δ T1 + m2c2 Δ T2

here m1 = 40 g

L1 = 80 cal/g ……. ice

C1 = 1

ΔT1 = 200 – 0 = 200 °C

m2 = 200 g

C2 = 1

ΔT2 = 100 – 0 = 100 °C

Heat required = 40 × 80 + 200 × 1 × 100 + 200 × 1 × 100 = 43200

Q = 43200 cal

Let ms is mass of steam

Q = msL2

for water steam L2 = 540 cal/g

ms = (Q / L2)

ms = [(43200) / (540)]

ms = 80 g – This is steam converted to water

Total water = 40 + 200 +80 = 320 g